(a) Assume the following information represents the National Income Model of an ‘Utopian’ economy.

Y = C + I + G
C = a + b(Y – T)
T = d + tY
I = IO
G = GO

Where a > O; O < b < 1
d > O; O < t < 1

T = Taxes
I = Investment
G = Government Expenditure

i) Explain the economic interpretation of the parameters a,b,d and t. (4 marks)
ii)Find the equilibrium values of income, consumption and taxes. (8 marks)
b) Discuss the three approaches used in measuring the national income of a country and show why they give the same estimate. (8 marks)
1
Expert's answer
2018-01-22T09:02:07-0500
i) a - autonomous consumption, that part that is consumed irrespective of the disposable income;
b - marginal propensity to consume;
d – autonomous level of taxes, that part that is paid irrespective of the disposable level of taxation;
t – tax rate.
ii) T = d + tY
C = a + b(Y-T) = a + b(Y-d-tY)
Y = C + I + G = a + bY – bd – btY + GO + IO
b) In the economy, the total product, total profit and total costs are always identical to each other. Thus, national income can be considered in terms of the triple identity between production, income and expenditure, that is,

National Product = National Income = National Expenses.

Consequently, there are three types of measuring system for national income. They are (i) the method of the product, (ii) the method of income and (iii) the method of expenditure.